PrettyEyes:

Isabelle is mixing red paint with blue paint to make purple paint. She adds of a fluid ounce of red to of a fluid ounce of blue to make fluid ounces of purple. How many fluid ounces of red paint will she need to make 3 fluid ounces of purple paint? A. fluid ounce B. fluid ounce C. fluid ounce D. fluid ounces

Angle:

Hi @PrettyEyes It looks like the fractions didn't copy over right - do you have a screenshot of the problem?

PrettyEyes:

no

Angle:

well your problem is missing a bunch of numbers...

PrettyEyes:

She adds 3/10 of a fluid ounce of red to 11/15 of a fluid ounce of blue to make 1 1/30 fluid ounces of purple.

Angle:

Those are some ugly numbers for a math problem... ugh ok... what is 3 divided by (1 1/30) = ?

Angle:

btw are your answer choices decimals or fractions?

PrettyEyes:

fractions

PrettyEyes:

it is 2 28/31

Angle:

awesome! now we need to do \(\huge \frac{3}{10}\times\large2\huge\frac{28}{31}\)

PrettyEyes:

27/31

Angle:

awesome, I think that should be your answer! :)

PrettyEyes:

it is thank you